A production facility contains two machines that are used to rework items that are initially defective. Let 𝑋 be the number of hours that the first machine is in use and let π‘Œ be the number of hours that the second machine is in use, on a randomly chosen day. Assume that 𝑋 and π‘Œ have a joint probability density function given by 𝑓(π‘₯) = { 3 2 (π‘₯ 2 + 𝑦 2 ) 0 < π‘₯ < 1 π‘Žπ‘›π‘‘ 0 < 𝑦 < 1 0 π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’. a. What is the probability that both machines are in operation for less than half an hour?

1
Expert's answer
2022-01-26T17:09:32-0500

To find the probability that both machines are in operation for less than half an hour, we determine the probability,

"p(0\lt x\lt 0.5,0\lt y\lt 0.5)=\displaystyle\int^{0.5}_0\displaystyle\int^{0.5}_0 f(x,y)dydx\\
=\displaystyle\int^{0.5}_0\displaystyle\int^{0.5}_0 {3\over2}(x^2+y^2)dydx\\
={3\over2}\displaystyle\int^{0.5}_0\displaystyle\int^{0.5}_0 (x^2+y^2)dydx
={3\over2}\displaystyle\int^{0.5}_0(x^2y+{y^3\over3})|^{0.5}_0dx\\
={3\over2}\displaystyle\int^{0.5}_0(0.5x^2+{1\over24})dx={3\over2}({1\over6}x^3+{x\over24})|^{0.5}_0={3\over2}({1\over48}+{1\over48})={1\over16}"

Therefore, the probability that both machines are in operation for less than half an hour is "{1\over16}"